# PALM: Pushing Adaptive Learning Rate Mechanisms for Continual Test-Time Adaptation

### v2

@@ 5.1 Experimental Settings and Baselines
BN-1 [28] 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
TENT Continual [36] 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
CoTTA [37]
@@
5.2 CTTA Classification Results
Results on CIFAR-10C and CIFAR-100C. We contrast our method against all the baselines as in Tables 1 and 2. The results reveal that our method performs the best across all training setups and baselines for both datasets. We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively. Concerning recent works of LAW and DSS, our method improves 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively. It is noteworthy that TENT Continual initially excels in the early tasks but experiences error accumulation with time, leading to an increase in mean error. In the study by CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)", the mean errors are better than most CTTA benchmarks, if not the best. LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the “domain-level" FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq. 1). This affirms the findings in [15] that concentrated adaptation, for image corruptions, in the initial layers suffices.
Results on Imagenet-C In Table 3, we report the CTTA results on the ImageNet-C dataset. As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)". We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively. In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
@@ 5.2 CTTA Classification Results
5.3 GTTA Classification Results
Results on CIFAR-10C, CIFAR-100C, and Imagenet-C. Here, we present the gradual test-time adaptation (GTTA) results on CIFAR-10C, CIFAR-100C, and Imagenet-C. For each task, the corruption severity of the inputs is gradually changed as discussed earlier. Table 4 reports the mean classification error for GTTA across all severities. We observe that, for all the datasets, our approach outperforms or performs comparably well across all the reported baselines. With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA, which also demonstrates competitive improvements in the mean classification error.
5.4 Ablation Studies
We conduct ablation experiments to validate our approach and gain insights into its underlying mechanisms and components. Specifically, we conduct these experiments and report results on the three CTTA benchmark datasets: CIFAR-10C, CIFAR-100C, and ImageNet-100C.
    WideResNet-28 [42]       ResNeXt-29 [39]       ResNet-50 [8]
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA. [Top row] - Variations for the domain “glass_blur" across datasets. [Bottom row] - Variations for the domain “snow".
@@ 5.4 Ablation Studies
Computational efficiency and Catastrophic forgetting: In our proposed method, we first select a subset of layers from the gradient space (Eq. 2) and adapt them while strictly freezing the unselected layers. In Table 5, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets. Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source model parameters. This showcases significant computational advantages over CoTTA, which updates 100% of the source parameters. Additionally, in LAW, the LRs of all the parameters are adjusted. It is worth mentioning that our method alleviates catastrophic forgetting (CF) by only adapting a small subset of model parameters ensuring that much of the prior source knowledge is preserved and CF is largely minimized.
@@
Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS [41], Prodigy [19], and Adafactor [32], for CTTA. We report the results in Table 6. The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
Analysis of LR importance $i_{j,n}^{t}$: The prime objective of our work is to adaptively alter the LRs of parameters of selected layers based on their sensitivity. In Figure 2, we illustrate how the LR importance $i_{j,n}^{t}$ (Eq. 7) of WideResNet-28, ResNeXt-29, and ResNet-50, averaged for a task, change during adaptation on their respective datasets. For brevity, we show two domains - glass_blur and snow. We observe that the importance varies across domains and models. Across all the models, we notice that generally, the deeper layers have smaller associated $i_{j,n}^{t}$, which means that the initial layers are adapted more to the target domain. This result completely aligns with the findings of surgical fine-tuning [15]. In addition to that, we also notice more sparsity in the deeper layers. This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
Layer importance via pseudo-labels: In LAW, the authors use the pseudo-labels to decide on the importance of layers for adaptation. In this, we study the effectiveness of such a design for our method i.e., we investigate if layer importance computed using pseudo-labels is helpful. We compute the gradients from the log-likelihood as $\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$ and estimate the sensitivity of the parameters of the chosen layers (Eq. 4-Eq. 6), and alter the learning rates (Eq. 7). Table 7 reports the mean classification errors for all the datasets. We observe that the gradients computed using the pseudo-labels are an unreliable indicator of prediction uncertainty. The scoring function $\mathcal{Z}_{\theta_{j,n}^{t}}=\|{{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}\|_{1}$ computed using these gradients carries insufficient information. We achieve improvements of 2.4%, 3.7%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
@@
a) Ablation on smoothing factor $\alpha$. b) Ablation on temperature T.
Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T for CIFAR-10C, CIFAR-100C, and ImageNet-C.
@@
Contribution of smoothing factor $\alpha$: A key component of our method is computing the moving average of sensitivities across data batches, smoothed by parameter $\alpha$ (Eq. 5). Fig. 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the “domain-level" sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes. This suggests that the effect of the current sensitivity of selected layers is indeed required. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW. For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
[…]

### v3

@@ 5 Experiments and Results
BN-1 (Schneider et al. 2020) 28.1 26.1 36.3 12.8 35.3 14.2 12.1 17.3 17.4 15.3 8.4 12.6 23.8 19.7 27.3 20.4
TENT Continual (Wang et al. 2020) 24.8 20.6 28.6 14.4 31.1 16.5 14.1 19.1 18.6 18.6 12.2 20.3 25.7 20.8 24.9 20.7
CoTTA (Wang et al. 2022)
@@
We evaluate our proposed method in a CTTA setting, comparing it to state-of-the-art approaches, and extend our study to the gradual test-time adaptation (GTTA) setting following (Wang et al. 2022; Park et al. 2024).
Experimental Settings and Baselines
Datasets and Tasks: Following the standard benchmarks set by (Wang et al. 2022), we evaluate our proposed method on CIFAR-10C, CIFAR-100C, and ImageNet-C, based on image corruption schemes as set in (Hendrycks and Dietterich 2019). Each dataset contains a set of 15 corruption styles as tasks (e.g. gaussian noise, shot noise, $\ldots$) with 5 severity levels. We follow CoTTA and apply the same test-time augmentations, like color jitter, gaussian blur, gaussian noise, random flips, and random affine, to the input $x_{k}$. As the evaluation metric, we report the mean classification error.
CTTA implementation details: To maintain fairness regarding the source models, we employ WideResNet-28 (Zagoruyko and Komodakis 2016) for CIFAR-10C, ResNeXt-29 (Xie et al. 2017) for CIFAR-100C, and ResNet-50 (He et al. 2016) for ImageNet-C, all available on RobustBench (Croce et al. 2020). We optimize using an Adam optimizer, setting base learning rates ($\kappa$) to 5e-4 for CIFAR-10C and CIFAR-100C, and 5e-5 for ImageNet-C. For balancing parameter sensitivities, we set $\alpha$ to 0.5, 0.9, and 0.5 respectively, with temperature coefficients T set to 50, 100, and 1000 respectively. We set $\eta$ to 1, 0.5, and 0.3, and $\lambda$ to 0.01 throughout. Batch sizes are set to 200, 200, and 64 for each dataset, following CoTTA, and results are only reported for the highest severity level of 5 for each task.
GTTA implementation details: In the standard CTTA setup, each task only receives images with the highest severity. However, in GTTA, following CoTTA, for each task, the severity levels are gradually changed as $\ldots$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\rightarrow$3$\rightarrow$4$\rightarrow$5$\rightarrow$4$\rightarrow$3$\rightarrow$2$\rightarrow$1 $change]{task}$ 1$\rightarrow$2$\ldots$. We adhere to the details specified in the CTTA setting.
Baselines: We compare against several CTTA baselines, including BN Stats Adapt (BN-1) (Schneider et al. 2020), TENT Continual (Wang et al. 2020), CoTTA (Wang et al. 2022), EcoTTA (Song et al. 2023), AdaContrast (Chen et al. 2022), DSS (Wang et al. 2024), and LAW (Park et al. 2024). We also contrast our work against surgical fine-tuning (Lee et al. 2022). For EcoTTA, we report the results on CIFAR-10C and ImageNet-C only since they do not use ResNeXt-29 (Xie et al. 2017) for CIFAR-100C. Following the findings of (Lee et al. 2022), we continually fine-tune only the first convolutional block of the respective models, denoted as “Surgical fine-tuning (continual)”, using an SGD optimizer with fixed learning rates of 1e-3 for the CIFAR datasets and 5e-4 for ImageNet-C. For the results of LAW, we reproduce their results strictly following their official implementation  [^1] https://github.com/junia3/LayerwiseTTA/tree/main with parameters as mentioned in their paper.
@@
CTTA Classification Results
Results on CIFAR-10C and CIFAR-100C: We report the CTTA results in Tables 1 and 2. Our method performs the best across all training setups and baselines for both datasets. We achieve large improvements of 4.9% and 5.2% for CIFAR-10C and 5.3% and 30.8% for CIFAR-100C, over BN-1 and TENT Continual respectively. Concerning recent works of LAW and DSS, our method improves by 1.7% and 0.5% on CIFAR-10C and 1.6% and 0.8% on CIFAR-100C, respectively. TENT Continual excels initially but accumulates errors over time, increasing mean error. In CoTTA, the method’s reliance on full model updates renders it computationally expensive and unsuitable for deployment. An interesting observation is that with a drastic reduction in the number of trainable parameters, as in “Surgical fine-tuning (continual)”, the mean errors are better than most CTTA benchmarks, if not the best. LAW uses the uncertain pseudo labels to calculate the domain shift approximated via the FIM. The minor error that is introduced for each batch begins to propagate across all the domains. In addition, the imbalance caused by the “domain-level” FIM and the current batch’s FIM, results in a higher error rate for the initial tasks. In our method, the computation of the model prediction uncertainty to automatically select layers is useful (Eq. 1). This affirms the findings in (Lee et al. 2022) that concentrated adaptation, for image corruptions, in the initial layers suffices.
Results on Imagenet-C: In Table 3, we report the CTTA results on the ImageNet-C dataset. As shown, our method outperforms all the previous CTTA baselines and “Surgical fine-tuning (continual)”. We observe an improvement of 0.9% and 4.5% over LAW and DSS respectively. In addition, we achieve 2.6%, 5.4%, and 2.5% improvements over CoTTA, AdaContrast, and TENT Continual, respectively.
@@
GTTA Classification Results
Results: Here, we present the GTTA results on the benchmark datasets. For each task, the corruption severity of the inputs is gradually changed as discussed earlier. Table 4 reports the mean classification error for GTTA across all severities. We observe that, for all the datasets, our approach outperforms or performs comparably well across all the reported baselines. With most parameters frozen during the adaptation on each batch, our method is more robust and inexpensive than CoTTA, which also demonstrates competitive improvements in the mean classification error.
Ablation Studies and Analysis
Computational efficiency and Catastrophic forgetting (CF): In our PALM, we first select a subset of layers from the gradient space (Eq. 2) and adapt them while strictly freezing the unselected layers. In Table 5, we report the average % of adapted model parameters per task with respect to the source model, for our method and CoTTA, across the benchmark datasets. Our method is computationally inexpensive and on an average updates 4.34%, 6.65%, and 17.21% of their respective source parameters. This showcases significant computational advantages over CoTTA, which updates 100% of the parameters. With a single NVIDIA A5000 GPU, PALM incurs a slightly higher adaptation time/batch, due to the two-stage proposed method. For example, on CIFAR-10, PALM runs for 0.8s per batch, whereas Surgical Fine-tuning (Surgical) and LAW take 0.45 s and 0.73 s respectively. Our method alleviates CF by only adapting a small subset of model parameters, ensuring that much of the prior source knowledge is preserved. Quantitatively, the source models achieved zero-shot mean accuracies of 94.03%, 76.06%, and 75.19% on respective datasets. After adaptation, PALM retained source knowledge better, with mean accuracies of 92.29%, 77.46%, and 70.87%, compared to LAW’s 91.67%, 77.69%, and 71.57%. Despite accuracy decline with more tasks, especially after adaptation on ImageNet-C, PALM matches or outperforms LAW.
@@
Adaptive LR optimizers vs Ours: Here, we study the effect of directly using adaptive LR optimizers like LARS (You, Gitman, and Ginsburg 2017), Prodigy (Mishchenko and Defazio 2023), and Adafactor (Shazeer and Stern 2018), for CTTA. We report the results in Table 6. The mentioned optimizers do not adjust layer-wise LRs based on parameter sensitivity, failing to capture rapid domain shifts, resulting in poorer performance compared to our method.
@@
    WideResNet-28     ResNeXt-29     ResNet-50
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 2: Illustrations of LR importance $i_{j,n}^{t}$ across different convolutional blocks/stages/layers during CTTA. [Top row] - Variations for the domain “glass_blur” across datasets. [Bottom row] - Variations for the domain “snow”.
Analysis of LR importance $i_{j,n}^{t}$: The prime objective of our work is to adaptively alter the LRs of parameters of selected layers based on their sensitivity. In Figure 2, we illustrate how the LR importance $i_{j,n}^{t}$ (Eq. 7) of WideResNet-28, ResNeXt-29, and ResNet-50, averaged for a task, change during adaptation on their respective datasets. For brevity, we show two domains - glass_blur and snow. We observe that the importance varies across domains and models. Across all the models, we notice that generally, the deeper layers have smaller associated $i_{j,n}^{t}$, which means that the initial layers are adapted more to the target domain. This result completely aligns with the findings of surgical fine-tuning (Lee et al. 2022). In addition to that, we also notice more sparsity in the deeper layers. This further confirms the necessity of freezing parameters automatically selected based on the target domain for adaptation.
Layer importance via pseudo-labels: LAW leverages pseudo-labels to determine layer importance for adaptation. We evaluate the effectiveness of this design for our method by investigating the utility of pseudo-label-based layer importance. We compute the gradients from the log-likelihood as $\nabla_{\theta_{j,n}^{t}}\log(p(x_{k}))$ and estimate the sensitivity of the parameters of the chosen layers (Eq. 4-Eq. 6), and alter the LRs (Eq. 7). Table 7 reports the mean classification errors for all the datasets. We observe that such gradients are unreliable indicators of prediction uncertainty. The scoring function $\mathcal{Z}_{\theta_{j,n}^{t}}=\|{{\nabla_{\theta_{j,n}^{t}}\mathcal{L}({\theta}^{t})}}\|_{1}$ computed using these gradients carries insufficient information. We achieve improvements of 2.4%, 3.7%, and 8.5% over CIFAR-10C, CIFAR-100C, and ImageNet-C, respectively.
@@
a) Ablation on $\alpha$. b) Ablation on temperature T.
Figure 3: Ablation results on smoothing factor $\alpha$ and temperature T.
@@
Contribution of smoothing factor $\alpha$: A key component of PALM is computing the moving average of sensitivities across data batches, smoothed by $\alpha$ (Eq. 5). Figure 3(a) illustrates the variations of the mean errors, for the benchmark datasets, for varying $\alpha$. We observe that the “domain-level” sensitivity across past batches of data is indeed useful. For $\alpha$ = 0 i.e., when $S_{j,n}^{t}$ of a selected parameter at timestep t is not considered, the respective datasets’ error rate drastically spikes. For CIFAR-10C, we achieve the best performance when $\alpha$ = 0.5 with an improvement of 1.7% over LAW. For CIFAR-100C and ImageNet-C, we see improvements of 1.6% and 0.9% over LAW, at $\alpha$ = 0.9 and 0.5 respectively. With higher values of $\alpha$, the results do not vary too much. Hence, attaining a balance between sensitivities from past domains ($\hat{S}_{j,n}^{t}$) and current sensitivity ($S_{j,n}^{t}$) is needed.
Effectiveness of temperature T: In our method, we compute the KL divergence between a uniform distribution u and a smoothed classifier’s output logits (Eq. 1), denoted by $\hat{h}(x_{k})$. To study its influence, we illustrate the results in Figure 3(b). Ideally, larger values of T push the distributions of the predictions closer to u (Hinton, Vinyals, and Dean 2015) by spreading the probability among inputs. In CTTA, since we deal with domain shifts with no labels, we argue that higher values of T are useful to capture the uncertainty. So, the KL divergence reduces, indicating more uncertainty in the model and hence, more adaptation. For a systematic temperature selection, we aim to adjust T based on the number of adapted parameters. In particular, if the number of adapted parameters increases across batches, T will be increased. For a difficult dataset like ImageNet-C, larger values of T are useful with the error reducing as T increases. For the CIFAR-C datasets, the errors begin to rise for T $\geq$ 100 which could be due to large spreads of the distributions.
@@ 6 Conclusion
Acknowledgments. We thank all the reviewers for their helpful comments and suggestions. This project was supported by a grant from the University of Texas at Dallas, Richardson, TX, USA.